OpenStudy (anonymous):

Why is this true? \[n=1: a_2=\frac{a_0}{2!}\] \[n=2: a_3=0\] \[n=3: a_4=\frac{a_0}{2^22!}\] \[n=4: a_5=0\] \[n=5: a_6=\frac{a_0}{2^32!}\] Why is it \[a_0\sum_{k=0}^{\infty} \frac{x^{2k}}{2^kk!}\] and not \[a_0\sum_{k=0}^{\infty} \frac{1}{2^kk!}\] isn't it: \[\frac{a_0}{2!}+\frac{a_0}{2^2(2!)}+\frac{a_0}{2^3(3!)}\] \[=a_0 \left(\frac{1}{2!}+\frac{1}{2^2(2!)}+\frac{1}{2^3(3!)}\]

OpenStudy (anonymous):

\[=a_0 \left(\frac{1}{2!}+\frac{1}{2^2(2!)}+\frac{1}{2^3(3!)}\right)\] error on n=5....should be: 2^3 (3!) in the denominator

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@Zarkon

OpenStudy (experimentx):

could you rectify that latex error

OpenStudy (anonymous):

yep it's in my second post the one with =a_0(....)

OpenStudy (experimentx):

 from the data you have given i could say ^^